# Interpreting Graph Neural Networks with In-Distributed Proxies

### v1

@@ 3.2 Derivation of Inner Optimization
For the inner optimization objective, we implement the first constraint by minimizing the distribution distance between $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$ and $P_{\mathcal{G}}$, under the Erdős–Rényi assumption, the distribution loss is equivalent to the cross-entropy loss between $\tilde{G}$ given $G^{\prime}$ and $G$ over the full adjacency matrix (Chen et al., 2023). Considering that $G$ is usually sparse compared to a fully connected graph, in practice, we adopt a weighted version to emphasize more connected node pairs (Wang et al., 2016). Formally, we have the following distribution loss.
@@
where $\bar{{\mathcal{E}}}$ is the set of node pairs that are unconnected in $G$, $\tilde{p}_{uv}$ is the probability of node pair $(u,v)$ in $\tilde{G}$, and $\beta$ is a hyper-parameter to get a trade-off between connected and unconnected node pairs.
The second constraint requires the mutual information of $Y$ and $\tilde{G}$ is the same as that of $Y$ and $G^{\prime}$. Due to the OOD problem, it is non-trivial to directly compute $P(Y|G^{\prime})$ or $H(Y|G^{\prime})$. Instead, we implement this constraint with a novel graph generator that $\tilde{G}$ is obtained by combining $G^{\prime}$ and a non-explanatory subgraph (Zhang et al., 2023b). In practice, we implement the non-explanatory part by perturbing the remaining subgraph $G-G^{\prime}$. The intuition is that if an explanation comprises information, it is hard to change the prediction by manipulating the remaining part that is non-explanatory, which is widely adopted in the literature (Fang et al., 2023a; Zhang et al., 2023b; Zheng et al., 2023).
Figure 2: ProxyExplainer consists of two components: an Explainer and a Proxy Graph Generator. The Explainer takes a graph $G$ as input and produces an explainable subgraph $G^{\prime}$, The proxy generator creates an in-distribution proxy graphs that preserve the label information in $G^{\prime}$. The proxy generator consists of a graph auto-encoder (GAE) and a variational graph auto-encoder (VGAE).
@@ 4 The ProxyExplainer
Inspired by our novel GIB with proxy graphs, in this section, we introduce a straightforward yet theoretically robust instantiation, named ProxyExplainer. As shown in Figure 2, the architecture of our model comprises two key modules: the explainer and the proxy graph generator. The explainer takes the $G$ as input and outputs a subgraph $G^{\prime}$ as an explanation, which is optimized through the outer objective in equation 9. The proxy graph generator generates an in-distributed proxy graph, which is optimized with the inner objective.
@@ 4.1 The Explainer
For the sake of optimization, we follow existing works to relax the element in ${\bm{A}}$ from binary values to continuous values in the range $[0,1]$ (Ying et al., 2019; Luo et al., 2020; Miao et al., 2022). We adopt a generative explainer due to its effectiveness and efficiency (Luo et al., 2020). Specifically, we decompose the to-be-explained model $f(\cdot)$ into two functions that $f_{\text{enc}}(\cdot)$ learns node representations and $f_{\text{cls}}(\cdot)$ predicts graph labels based on node embeddings. Formally, we have ${\bm{Z}}=f_{\text{enc}}({\bm{X}},{\bm{A}})$ and $Y=f_{\text{cls}}({\bm{Z}})$, where ${\bm{Z}}\in{\mathbb{R}}^{N\times F}$ is the node representation matrix and $F$ is the number of feature dimenions. Routinely, $f_{\text{cls}}(\cdot)$ consists of a pooling layer followed by a classification layer and $f_{\text{enc}}(\cdot)$ consists of the other layers. Following the independence assumption in Section 3.1, we approximate Bernoulli distributions with binary concrete distributions (Maddison et al., 2017; Jang et al., 2017). Specifically, the probability of sampling an edge $(i,j)$ is computed by an MLP parameterized by $\psi$, denoted by $g_{\psi}(\cdot)$. Formally, we have
$\omega_{ij}=g_{\psi}([{\bm{z}}_{i};{\bm{z}}_{j}]),\quad\quad\epsilon\sim\text{Uniform}(0,1)$ (13)
${A}_{ij}^{\prime}=\sigma((\log\epsilon-\log(1-\epsilon)+\omega_{ij})/\tau),$
where $[{\bm{z}}_{i};{\bm{z}}_{j}]$ is the concatenation of node representations ${\bm{z}}_{i}$ and ${\bm{z}}_{j}$, $\epsilon$ is a independent variable, $\sigma(\cdot)$ is the Sigmoid function, and $\tau$ is a temperature hyper-parameter for approximation.
@@ 4.2 The Proxy Graph Generator
Table 1: Explanation accuracy in terms of AUC-ROC on edges under six datasets based on GCN.
BA-2motifs
BA-3motifs
Alkane-Carbonyl
Benzene
Fluoride-Carbonyl
MUTAG
GradCAM
0.714±${}_{\pm 0.000}$
0.709±${}_{\pm 0.000}$
0.496±${}_{\pm 0.000}$
0.662±${}_{\pm 0.000}$
0.604±${}_{\pm 0.000}$
0.573±${}_{\pm 0.000}$
GNNExplainer
0.644±${}_{\pm 0.007}$
0.511±${}_{\pm 0.002}$
0.532±${}_{\pm 0.008}$
0.499±${}_{\pm 0.004}$
0.540±${}_{\pm 0.002}$
0.682±${}_{\pm 0.009}$
PGExplainer
0.734±${}_{\pm 0.117}$
0.796±${}_{\pm 0.010}$
0.611±${}_{\pm 0.071}$
0.698±${}_{\pm 0.099}$
0.591±${}_{\pm 0.041}$
0.832±${}_{\pm 0.032}$
ReFine
0.698±${}_{\pm 0.001}$
0.629±${}_{\pm 0.005}$
0.630±${}_{\pm 0.007}$
0.533±${}_{\pm 0.011}$
0.507±${}_{\pm 0.003}$
0.612±${}_{\pm 0.004}$
MixupExplainer
0.906±${}_{\pm 0.059}$
0.859±${}_{\pm 0.019}$
0.642±${}_{\pm 0.033}$
0.520±${}_{\pm 0.043}$
0.534±${}_{\pm 0.012}$
0.883±${}_{\pm 0.103}$
ProxyExplainer
0.935±${}_{\pm 0.008}$ 0.960±${}_{\pm 0.008}$ 0.806±${}_{\pm 0.003}$ 0.745±${}_{\pm 0.012}$ 0.641±${}_{\pm 0.070}$ 0.977±${}_{\pm 0.009}$
@@
where $\mathcal{L}_{\text{dist}}$ is the distribution distance between $\tilde{G}$ and $G$. $\mathcal{L}_{\text{KL}}$ represents the Kullback-Leibler (KL) divergence between the distribution of the latent representations ${\bm{Z}}^{\Delta}$ and the assumed Gaussian prior. This term is crucial for regulating the variational aspect of the VGAE, ensuring that the generated perturbations are meaningful and controlled. $\lambda$ is the hyper-parameter.
Alternative Training. To train explainer and proxy graph generator networks, we follow existing works (Zheng et al., 2023) to use an alternating training schedule that trains the proxy graph generator network $M$ times and then trains the explainer network one time. $M$ is a hyper-parameter determined by grid search. The detailed algorithm description of our model is shown in Appendix B.
@@ 5 Related Work
GNN Explanation. The goal of explainability in GNNs is to ensure transparency in graph-based tasks. Recent works have been directed towards elucidating the rationale behind GNN predictions. These explanation methods can be broadly classified into two categories: instance-level and model-level approaches (Yuan et al., 2022). In this study, we focus on instance-level explanations, which aim to clarify the specific reasoning behind individual predictions made by GNNs. These methods are critical for understanding the decision-making process on a case-by-case basis to enhance the explainability of GNNs in graph classification tasks. For example, GNNExplainer (Ying et al., 2019) excludes certain edges and node features to observe the changes in classification. However, its single-instance focus limits its applicability to provide a global understanding of the to-be-explained model. PGExplainer (Luo et al., 2020) introduces a parametric neural network to learn edge weights. Thus, once training is complete, it can explain new graphs without retraining. ReFine (Wang et al., 2021) integrates a pre-training phase that focuses on class comparisons and is fine-tuned to refine context-specific explanations. GStarX (Zhang et al., 2022b) assigns an importance score to each node by caculating the Hamiache and Navarro values of the structure to obtain explanatory subgraphs. GFlowExplainer (Li et al., 2023) uses a generator to construct a TD-like flow matching condition to learn a policy for generating explanations by adding nodes sequentially.
Distribution Shifting in Explanations. The distribution shifting problem in post-hoc explanations has been increasingly recognized in explainable AI fields (Chang et al., 2019; Qiu et al., 2022). For example, FIDO (Chang et al., 2019) works on enhancing image classifier explanations, focusing on relevant contextual details that agree with the training data’s distribution. A recent study tackles the distribution shifting problem in image explanations by introducing a module that assesses the similarity between altered data and the original dataset distribution (Qiu et al., 2022). In the graph domain, an ad-hoc strategy to mitigate distribution shifting is to initially reduce the size constraint coefficient during the explanation process (Fang et al., 2023c). MixupExplainer (Zhang et al., 2023b) proposes a non-parametric solution by mixing up the explanation subgraph with a non-explainable part from another graph. However, this method operates under the assumption that the explanatory and non-explanatory subgraphs in mixed graphs are independent, which may not hold in many real-life graphs.
@@ 6 Experiments
Figure 3: Ablation Studies on BA-2motifs and Alkane-Carbonyl.
@@ 6.1 Experimental Settings
To evaluate the performance of ProxyExplainer, we use six benchmark datasets with ground-truth explanations. These include two synthetic datasets: BA-2motifs (Luo et al., 2020) and BA-3motifs (Chen et al., 2023), along with four real-world datasets: Alkane-Carbonyl (Sanchez-Lengeling et al., 2020), Benzene (Sanchez-Lengeling et al., 2020), Fluoride-Carbonyl (Sanchez-Lengeling et al., 2020), and MUTAG (Kazius et al., 2005). We take GradCAM (Pope et al., 2019), GNNExplainer(Ying et al., 2019), PGExplainer(Luo et al., 2020), ReFine (Wang et al., 2021), and MixupExplainer (Zhang et al., 2023b) for comparison. We follow the experimental setting in previous works (Ying et al., 2019; Luo et al., 2020; Sanchez-Lengeling et al., 2020) to train a Graph Convolutional Network (GCN) model with three layers. We use the Adam optimizer (Kingma & Ba, 2014) with the inclusion of a weight decay $45e-$ 4. To evaluate the quality of explanations, we approach the explanation task as a binary classification of edges. Edges that are part of ground truth subgraphs are labeled as positive, while all others are deemed negative. The importance weights given by the explanation methods are interpreted as prediction scores. An effective explanation technique is one that assigns higher weights to edges within the ground truth subgraphs compared to those outside of them. For quantitative assessment, we utilize the AUC-ROC metric. Detailed information regarding datasets and baselines is delineated in the Appendix C.

### v2

@@ 3.2 Derivation of Inner Optimization
For the inner optimization objective, we implement the first constraint by minimizing the distribution distance between $Q_{\bm{\phi}}(\tilde{G}|G^{\prime})$ and $P_{\mathcal{G}}$. Under the Erdős–Rényi assumption, the distribution loss is equivalent to the cross-entropy loss between $\tilde{G}$ given $G^{\prime}$ and $G$ over the full adjacency matrix (Chen et al., 2023b). Considering that $G$ is usually sparse rather than a fully connected graph, in practice, we adopt a weighted version to emphasize more connected node pairs (Wang et al., 2016). Formally, we have the following distribution loss.
@@
where $\bar{{\mathcal{E}}}$ is the set of node pairs that are unconnected in $G$, $\tilde{p}_{uv}$ is the probability of node pair $(u,v)$ in $\tilde{G}$, and $\beta$ is a hyper-parameter to for the trade-off between connected and unconnected node pairs.
The second constraint requires the mutual information of $Y$ and $\tilde{G}$ is the same as that of $Y$ and $G^{\prime}$. Due to the OOD problem, it is non-trivial to directly compute $P(Y|G^{\prime})$ or $H(Y|G^{\prime})$. Instead, we implement this constraint with a novel graph generator that $\tilde{G}$ is obtained by combining $G^{\prime}$ and a non-explanatory subgraph (Zhang et al., 2023b). In practice, we implement the non-explanatory part by perturbing the residual subgraph $G-G^{\prime}$. The intuition is that if an explanation comprises label information, it is unlikely to change the prediction by manipulating the remaining non-explanatory part, which is widely adopted in the literature (Fang et al., 2023a; Zheng et al., 2024).
Figure 2: ProxyExplainer consists of two components: an Explainer and a Proxy Graph Generator. The Explainer takes a graph $G$ as input and produces an explainable subgraph $G^{\prime}$, The proxy generator creates in-distribution proxy graphs that preserve the label information in $G^{\prime}$. The proxy generator consists of a graph auto-encoder (GAE) and a variational graph auto-encoder (VGAE).
@@ 4 The ProxyExplainer
Based on our novel GIB with proxy graphs, in this section, we introduce a straightforward yet theoretically robust instantiation, named ProxyExplainer. As shown in Figure 2, the architecture of our model comprises two key modules: the explainer and the proxy graph generator. The explainer takes the original graph $G$ as its input and outputs a subgraph $G^{\prime}$ as an explanation, which is optimized through the outer objective in equation 9. The proxy graph generator generates an in-distributed proxy graph, which is optimized with the inner objective.
@@ 4.1 The Explainer
For the sake of optimization, we follow the existing works to relax the element in ${\bm{A}}$ from binary values to continuous values in the range $[0,1]$ (Luo et al., 2020, 2024). We adopt a generative explainer due to its effectiveness and efficiency (Chen et al., 2023a). Specifically, we decompose the to-be-explained model $f(\cdot)$ into two functions that $f_{\text{enc}}(\cdot)$ learns node representations and $f_{\text{cls}}(\cdot)$ predicts graph labels based on node embeddings. Formally, we have ${\bm{Z}}=f_{\text{enc}}({\bm{X}},{\bm{A}})$ and $\hat{Y}=f_{\text{cls}}({\bm{Z}})$, where ${\bm{Z}}$ is the node representation matrix and $\hat{Y}$ is prediction. Routinely, $f_{\text{cls}}(\cdot)$ consists of a pooling layer followed by a classification layer and $f_{\text{enc}}(\cdot)$ consists of the other layers. Following the independence assumption in Section 3.1, we approximate Bernoulli distributions with binary concrete distributions (Maddison et al., 2017; Jang et al., 2017). Specifically, the probability of sampling an edge $(u,v)$ is computed by an MLP parameterized by $\psi$, denoted by $g_{\psi}(\cdot)$. Formally, we have
$\omega_{uv}=g_{\psi}([{\bm{z}}_{u};{\bm{z}}_{v}]),$ (13)
$\epsilon\sim\text{Uniform}(0,1),$
${A}_{uv}^{\prime}=\sigma((\log\epsilon-\log(1-\epsilon)+\omega_{uv})/\tau),$
where $[{\bm{z}}_{u};{\bm{z}}_{v}]$ is the concatenation of node representations ${\bm{z}}_{u}$ and ${\bm{z}}_{v}$, $\epsilon$ is a independent variable, $\sigma(\cdot)$ is the Sigmoid function, and $\tau$ is a temperature hyper-parameter for approximation. According to (Luo et al., 2020), the parameter $\pi^{\prime}_{uv}$ of Bernoulli distribution in equation 11 can be obtained by $\pi^{\prime}_{uv}=\frac{\exp(\omega_{uv})}{1+\exp(\omega_{uv})}$.
@@ 4.2 The Proxy Graph Generator
where $\mathcal{L}_{\text{dist}}$ is equivalent to cross-entropy between $\tilde{G}$ and $G$ (Chen et al., 2023b). $\mathcal{L}_{\text{KL}}$ represents the KL divergence between the distribution of the latent representations ${\bm{Z}}^{\Delta}$ and the assumed Gaussian prior. This term is crucial for regulating the variational aspect of the VGAE, ensuring that the generated perturbations are meaningful and controlled. $\lambda$ is a hyper-parameter.
Alternate Training. To train the explainer and the proxy graph generator networks, we follow existing works (Zheng et al., 2024) to use an alternate training schedule that trains the proxy graph generator network $M$ times and then trains the explainer network once. $M$ is a hyper-parameter determined by grid search. The detailed algorithm description of our model is shown in Appendix B.
@@ 5 Related Work
GNN Explanation. The goal of explainability in GNNs is to ensure transparency in graph-based tasks. Recent works have been directed towards elucidating the rationale behind GNN predictions. These explanation methods can be broadly classified into two categories: instance-level and model-level approaches (Yuan et al., 2022). In this study, we focus on instance-level explanations, which aim to clarify the specific reasoning behind individual predictions made by GNNs. These methods are critical for understanding the decision-making process on a case-by-case basis to enhance the explainability of GNNs. For example, GNNExplainer (Ying et al., 2019) excludes certain edges and node features to observe the changes in prediction. However, its single-instance focus limits its applicability to provide a global understanding of the to-be-explained model (Chen et al., 2023a). PGExplainer (Luo et al., 2020, 2024) introduces a parametric neural network to learn edge weights. Thus, once training is complete, it can explain new graphs without retraining. ReFine (Wang et al., 2021) integrates a pre-training phase that focuses on class comparisons and is fine-tuned to refine context-specific explanations. GStarX (Zhang et al., 2022) assigns an importance score to each node by caculating the Hamiache and Navarro values of the structure to obtain explanatory subgraphs. GFlowExplainer (Li et al., 2023) uses a generator to construct a TD-like flow matching condition to learn a policy for generating explanations by adding nodes sequentially.
Distribution Shifting in Explanations. The distribution shifting problem in post-hoc explanations has been increasingly recognized in explainable AI fields (Chang et al., 2019; Qiu et al., 2022). For example, FIDO (Chang et al., 2019) works on enhancing image classifier explanations, focusing on relevant contextual details that agree with the training data’s distribution. A recent study tackles the distribution shifting problem in image explanations by introducing a module that assesses the similarity between altered data and the original dataset distribution (Qiu et al., 2022). In the graph domain, an ad-hoc strategy to mitigate distribution shifting is to initially reduce the size constraint coefficient during the explanation process (Fang et al., 2023b). MixupExplainer (Zhang et al., 2023b) and RegExplainer (Zhang et al., 2023a) propose non-parametric solutions by mixing up the explanation subgraph with a non-explainable part from another graph. However, these methods operates under the assumption that the explanatory and non-explanatory subgraphs in mixed graphs are independent, which may not hold in many real-life graphs.
@@ 6 Experiments
Table 1: Explanation accuracy in terms of AUC-ROC on edges.
MUTAG Benzene Alkane-Carbonyl Fluoride-Carbonyl BA-2motifs BA-3motifs
GradCAM 0.727±0.000 0.740±0.000 0.448±0.000 0.694±0.000 0.714±0.000 0.709±0.000
GNNExplainer 0.682±0.009 0.485±0.001 0.551±0.003 0.574±0.002 0.644±0.007 0.511±0.002
PGExplainer 0.832±0.032 0.793±0.054 0.660±0.036 0.702±0.018 0.734±0.117 0.796±0.010
ReFine 0.612±0.004 0.606±0.002 0.768±0.001 0.571±0.000 0.698±0.001 0.629±0.005
MixupExplainer 0.863±0.103 0.611±0.032 0.811±0.006 0.706±0.013 0.906±0.059 0.859±0.019
ProxyExplainer 0.977±0.009 0.845±0.036 0.934±0.005 0.758±0.068 0.935±0.008 0.960±0.008
Table 2: Explanation accuracy in terms of AP on MUTAG and BA-2motifs.
MUTAG BA-2motifs
GradCAM $0.247_{\pm 0.000}$ $0.664_{\pm 0.000}$
GNNExplainer $0.232_{\pm 0.001}$ $0.608_{\pm 0.004}$
PGExplainer $0.611_{\pm 0.024}$ $0.682_{\pm 0.117}$
ReFine $0.227_{\pm 0.001}$ $0.619_{\pm 0.002}$
MixupExplainer $0.647_{\pm 0.083}$ $0.787_{\pm 0.073}$
ProxyExplainer $\textbf{0.756}_{\pm 0.107}$ $\textbf{0.839}_{\pm 0.036}$
Table 3: Fidelity evaluation on MUTAG and BA-2motifs.
MUTAG BA-2motifs
$Fid_{\alpha_{1},{+}}\uparrow$ $Fid_{\alpha_{2},{-}}\downarrow$ $Fid_{\alpha_{1},{+}}\uparrow$ $Fid_{\alpha_{2},{-}}\downarrow$
GradCAM 0.004±0.000 $0.162_{\pm 0.000}$ $0.072_{\pm 0.000}$ $0.107_{\pm 0.000}$
GNNExp. $0.031_{\pm 0.001}$ $0.148_{\pm 0.001}$ $0.057_{\pm 0.002}$ $0.132_{\pm 0.001}$
PGExp. $0.034_{\pm 0.011}$ $0.148_{\pm 0.005}$ $0.065_{\pm 0.017}$ $0.126_{\pm 0.009}$
ReFine $0.003_{\pm 0.000}$ $0.160_{\pm 0.001}$ $0.060_{\pm 0.005}$ $0.125_{\pm 0.001}$
MixupExp. $0.037_{\pm 0.006}$ $0.146_{\pm 0.003}$ $0.074_{\pm 0.005}$ $0.112_{\pm 0.003}$
ProxyExp. $\textbf{0.040}_{\pm 0.002}$ $\textbf{0.145}_{\pm 0.001}$ $\textbf{0.086}_{\pm 0.003}$ $\textbf{0.106}_{\pm 0.002}$
@@ 6.1 Experimental Settings
To evaluate the performance of ProxyExplainer, we use six benchmark datasets with ground-truth explanations. These include four real-world datasets: MUTAG (Kazius et al., 2005), Benzene (Sanchez-Lengeling et al., 2020),Alkane-Carbonyl (Sanchez-Lengeling et al., 2020), and Fluoride-Carbonyl (Sanchez-Lengeling et al., 2020), along with two synthetic datasets: BA-2motifs (Luo et al., 2020) and BA-3motifs (Chen et al., 2023b). We take GradCAM (Pope et al., 2019), GNNExplainer(Ying et al., 2019), PGExplainer(Luo et al., 2020), ReFine (Wang et al., 2021), and MixupExplainer (Zhang et al., 2023b) for comparison. We follow the experimental setting in previous works (Ying et al., 2019; Luo et al., 2020; Sanchez-Lengeling et al., 2020) to train a Graph Convolutional Network (GCN) model (Kipf & Welling, 2017) with three layers. Experiments on another representative GNN, Graph Isomorphism Network (GIN) (Xu et al., 2019), can be found in Appendix D.3. We use the Adam optimizer (Kingma & Ba, 2014) with the inclusion of a weight decay $45e-$ 4. Detailed information regarding datasets and baselines is delineated in Appendix C.
To evaluate the quality of explanations, we approach the explanation task as a binary classification of edges. Edges that are part of ground truth subgraphs are labeled as positive, while all others are deemed negative. The importance weights given by the explanation methods are interpreted as prediction scores. An effective explanation technique is one that assigns higher weights to edges within the ground truth subgraphs compared to those outside of them. We utilize the AUC-ROC for quantitative assessment (Ying et al., 2019; Luo et al., 2020).
@@ 6.2 Quantitative Evaluation (RQ1)
The results demonstrate that ProxyExplainer provides the most accurate explanations across all datasets. Specifically, it improves the AUC scores by an average of $10.6\%$ on real-world datasets and $7.5\%$ on synthetic datasets over the leading baselines. Comparisons with baseline methods highlight the advantages of our proposed explanation framework. Besides, ProxyExplainer captures underlying explanatory factors consistently across diverse datasets. For instance, MixupExplainer exhibits proficiency on the synthetic BA-2motifs dataset but performs poorly on the real-world Benzene dataset. The reason is that MixupExplainer relies on the independence assumption of explanation and non-explanation subgraphs, which may not hold in real-world datasets. In contrast, ProxyExplainer consistently demonstrates high performance across different datasets, showcasing its robustness and adaptability.
Considering the importance of precision for the positive class in our context, we further adopt AP to evaluate the performance. As shown in Table 2, with AP scores, ProxyExplainer consistently outperforms the other baselines in two benchmark datasets MUTAG and BA-2motifs.
Additionally, some existing works, such as SubgraphX (Yuan et al., 2021), adopt faithfulness-based metrics for evaluation. However, these metrics are problematic due to the OOD problem (Zheng et al., 2024; Amara et al., 2023). So we use the robust fidelity metrics ($Fid_{\alpha_{1},{+}}$, $Fid_{\alpha_{2},{-}}$) as described in (Zheng et al., 2024) with default settings ($\alpha_{1}=0.1,\alpha_{2}=0.9$) to evaluate model faithfulness. Table 3 demonstrates that ProxyExplainer is consistently outperforms all baselines in both $Fid_{\alpha_{1},{+}}$ and $Fid_{\alpha_{2},{-}}$.
6.3 Alleviating Distribution Shifts (RQ2)
In this section, we assess ProxyExplainer’s ability to generate in-distribution proxy graphs. Due to the intractable of direct computation, we follow the previous work (Chen et al., 2023b) to utilize Maximum Mean Discrepancy (MMD) between distributions of multiple graph statistics, including degree distributions, clustering coefficients, and spectrum distributions, between the generated proxy graphs and original graphs. Specifically, we utilize Gaussian Earth Mover’s Distance kernel when computing MMDs. Smaller MMD values indicate similar graph distributions. For comparison, we also include the Ground truth explanations and the ones generated by PGExplainer.
MUTAG Benzene Alkane-Carbonyl
Metric GT PGE Proxy GT PGE Proxy GT PGE Proxy
Deg. 0.614 0.468 0.123 0.843 0.393 0.236 0.872 0.665 0.177
Clus. 0.003 0.003 0.009 0.009 0.002 0.004 0.011 0.011 0.011
Spec. 0.414 0.341 0.186 0.295 0.163 0.101 0.596 0.447 0.049
Sum. 1.032 0.813 0.317 1.147 0.558 0.341 1.479 1.123 0.237
Fluoride-Carbonyl BA-2motifs BA-3motifs
Metric GT PGE Proxy GT PGE Proxy GT PGE Proxy
Deg. 0.638 0.488 0.196 0.759 0.496 0.060 0.541 0.149 0.092
Clus. 0.012 0.012 0.012 0.447 0.463 0.584 0.262 0.382 0.245
Spec. 0.351 0.315 0.100 0.245 0.256 0.091 0.217 0.063 0.062
Sum. 1.000 0.815 0.308 1.451 1.215 0.735 1.020 0.594 0.399
The results are shown in Table 4. “GT” denotes the MMDs between the ground truth explanations and original graphs. “PGE” represents the MMDs between explanations generated by PGExplainer and original graphs. “Proxy” denotes the MMDs between proxy graphs in our method and original graphs. We have the following observations. First, the MMDs between ground truth and original graphs are usually large, verifying our motivation that a model trained on original graphs may not have correct predictions on the OOD explanation subgraphs. Second, the explanations generated by a representative work, PGExplainer, are often OOD from original graphs, indicating that the original GIB-based objective function may be sub-optimal. Third, in most cases, proxy graphs generated by our method are with smaller MMDs, demonstrating their in-distribution property.
6.4 Ablation Studies (RQ3)
In this section, we conduct ablation studies to investigate the roles of different components. Specifically, we consider the following variants of ProxyExplainer: (1) w/o $G^{\Delta}$: in this variant, we remove the non-explanatory subgraph generator (VGAE), which is the bottom half as shown in Figure 2; (2) w/o $\mathcal{L}_{\text{KL}}$: in this variant, we remove the KL divergence from the training loss in ProxyExplainer; (3) w/o $\mathcal{L}_{\text{dist}}$: in this variant, we remove the distribution loss from ProxyExplainer. The results of the ablation study on MUTAG and BA-2motifs are reported in Figure 3.
Figure 3 illustrates a notable performance drop for all variants, indicating that each component contributes positively to the effectiveness of ProxyExplainer. Especially, in the real-life dataset MUTAG, without the in-distribution constraint, w/o $\mathcal{L}_{\text{dist}}$ is much worse than ProxyExplainer, indicating the vital role of in-distributed proxy graphs in our framework. Extensive ablation studies on other datasets can be found in the Appendix D.2.
Figure 3: Ablation Studies on MUTAG and BA-2motifs.